# Unlocking Open-Set Language Accessibility in Vision Models

### v1

@@ Abstract
Visual classifiers offer high-dimensional feature representations that are challenging to interpret and analyze. Text, in contrast, provides a more expressive and human-friendly interpretable medium for understanding and analyzing model behavior. We propose a simple, yet powerful method for reformulating any visual classifier so that it can be accessed with open-set text queries without compromising its original performance. Our approach is label-free, efficient, and preserves the underlying classifier’s distribution and reasoning processes. We thus unlock several text-based interpretability applications for any classifier. We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language. In both applications, we achieve state-of-the-art results, greatly outperforming existing works. Our method enables text approaches for interpreting visual classifiers.
Machine Learning, ICML
@@ 1 Introduction
Visual classifiers inherently and purely provide visual signals, which are high-dimensional dense vector representations of images that are hard to interpret and dissect, limiting attribution maps as the main medium for interpretation. Attribution maps provide general and high-level information, indicating where in the image the most important region is for the prediction. Nevertheless, text is significantly more expressive, human-friendly, specific, and unlocks a variety of other interpretability applications that require language (Tewel et al., 2021; Koh et al., 2020; Oikarinen & Weng, 2023; Song et al., 2022; Minderer et al., 2022). Therefore, a question that arises is how can we talk or access those visual classifiers via open-set text? Performing this allows us to have a text interface that can be used to query visual classifiers. So far, this has only been achieved with generic vision-language models such as CLIP-based ones (Radford et al., 2021), which learn a shared vision-language space and thus enables any text to query the visual model directly. However, the deep learning field extends far beyond CLIP. For example, industries—being the primary end-users of such models—often rely on custom visual models tailored to their specific needs. This creates an application bottleneck: any task involving both visual and textual interpretations becomes restricted to CLIP or similar approaches, limiting flexibility, forcing reliance on such models, and inheriting their limitations.
Figure 1: We formulate any classifier to allow querying it and measuring its association with any text, while maintaining its performance and distribution. $f$ is the feature vector of an image.
@@
In this work, we lift this restriction by proposing a method to reformulate any visual classifier such that it can be accessed via open-set text (Figure 1). Our method is characterized by four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, it is faithful to the classifier and preserves its underlying distribution and reasoning process. Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
@@
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. In ImageNet-1K (Deng et al., 2009), there are a total of $1,000$ labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are embedded into vector representations (e.g, using a text encoder or word embedding model), they provide semantic information. Specifically, these embeddings also semantically encode words that are close to the class name within the text embedding space. For the “goldfish” example, the closest words could be “freshwater”, “fins” and “orange”. Our method learns to map images into this text embedding space, thus associating both the class name and its surrounding semantic space with the image. This is accomplished through a trainable MLP that projects the visual features into the text embedding space, and is trained by aligning its output distribution across all classes with the original classifier’s output distribution, while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with any text query beyond the class names.
@@
We demonstrate the effectiveness of our method with two primary applications. 1) We build label-free concept bottleneck models for any classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the new linear classifier remains faithful to the original classifier. 2) We show that our method can decode visual features of any classifier into natural language in a zero-shot manner, without requiring training on image-text data such as image-caption pairs, while also showing the generalization of our method to datasets beyond ImageNet.
In summary, our contributions are as follows:
  • We propose an efficient, label-free method to reconfigure any visual classifier such that it can be queried with open-set text queries, while also being faithful to the classifier.
  • We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free, faithful zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language.
  • Our method sets new state-of-the-art results outperforming existing works, including those which use CLIP as the visual encoder, although being trained with 400$\times$ less images and 400,000$\times$ less captions, thus efficiently learning faithful text interfaces to interpret any visual classifier.
@@ 2 Related Work
While CLIP is the predominant approach to interpret vision models through language (Oikarinen & Weng, 2023), there exist works that try to decode visual features of models beyond CLIP. DeVIL (Dani et al., 2023) trains an autoregressive text generator to map visual features from different layers of a classifier into image captions, leveraging annotated image-caption pairs as ground-truth data. Similarly, Natural Language Explanations (NLEs) (Park et al., 2018; Kayser et al., 2021; Sammani et al., 2022) use annotated textual explanations in place of conventional captions. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, and are therefore not faithful to the classifier. ZS-A2T (Salewski et al., 2023) maps attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective), and thus cannot be applied to visual classifiers. Our work on the other hand decodes visual features of any classifier in a zero-shot manner without requiring any annotated data, while also maintaining faithfulness to the original classifier.
Text-to-Concept (T2C) (Moayeri et al., 2023) is the closest related work to ours, where a linear layer is trained to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Since the linear layer maps features to the CLIP space, T2C is strongly biased towards interpreting the CLIP model rather than the original classifier. As the ground-truth data for training the linear mapper are the features extracted from the CLIP vision encoder, this method is approximate to using the CLIP vision encoder directly to encode the image. Moreover, T2C relies on CLIP supervision. In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution, ensuring faithfulness to it.
@@ 3 Method
Figure 2: An overview of our method for reformulating visual classifiers. (a) The process of training the MLP mapping between vision and text space. (b) The process of inference with adapted visual classifier. The text encoder acts as a linear classifier weight generator in our re-formulated classification process. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
@@
We propose to learn a simple light-weight MLP mapping function which projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$. Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method efficient. We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}” where {class} represents the class name in text format. This results in $K$ textual prompts, each of which is encoded with $T$: $u_{i}=T(l^{p}_{i})$, $\forall i=1,\dots,K$. Stacking all the encoded prompts, we get a matrix $U\in\mathbb{R}^{K\times m}$. We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$. Equivalently, this can be performed as a single matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class.
The most straightforward approach to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions). Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original distribution $o$ through the cross-entropy loss. For a single sample, the loss is given by:
@@
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but distill the distribution of the original model to a reformulated way of classification. This makes our approach label-free and preserves the distribution and reasoning process of the classifier $F$. We provide a PyTorch-like pseudocode of our training approach in Listing LABEL:lst:pseudo, and illustrate it in Figure 2(a). It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions or hierarchies (see Section F in the Appendix for more information).
After training, the projected visual features and the text encoder features lie in the same space, and we can therefore query the visual features with any text by finding the alignment score between the encoded text and the visual features. In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the newly formulated classification task, defined as argmax($\tilde{f}.U^{T}$). This is shown in Figure 2(b). We show later that this weight generating process is important to build faithful concept bottleneck models.
# mlp: trainable MLP from visual_dim -> text_dim
# images: batch of B images, shape (N, 3, height, width)
visual_feats = vision_encoder(images) # (N, visual_dim)
logits = visual_feats @ classifier # (N, num_classes)
original_dist = softmax(logits, dim=-1) # (N, num_classes)
mapped_feats = mlp(visual_feats) # (N, text_dim)
mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
text_feats = l2_norm(text_feats) # (N, text_dim)
pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
# cross entropy with original model’s soft distribution
loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
loss.backward() # only mapper parameters are updated
Listing 1: PyTorch-like pseudocode for our method
@@ 4.1 Zero-Shot Concept Bottleneck Models
Concept Bottleneck Models (CBMs) (Koh et al., 2020) are a class of inherently interpretable models and have recently attracted significant attention. They consist of two steps: (1) concept discovery, followed by (2) concepts-to-predictions. In (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score which represents the concept activation to the image. In (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of dense features. Recently, the Label-Free CBMs (LF-CBMs) family uses CLIP to perform the concept discovery step without annotated image-concept data, by either using CLIP as a model to provide ground-truth image-concept similarities to train a layer which performs concept discovery (Oikarinen et al., 2023), or by querying the image features from a set of predefined concepts within the CLIP space (Yang et al., 2022; Panousis et al., 2023) and using the cosine similarity scores between the concepts and the image as concept activations. However, there are major issues in the current literature of CBMs. Firstly, LF-CBMs are restricted to CLIP-based models, as the concept discovery step would otherwise not be possible since it necessitates an image-text similarity model. Secondly, all CBM approaches to date require training a linear probe on the concept activations which 1) hinders an on-the-fly ready-to-deploy CBMs, and 2) disconnects $W^{con}$ from the original feature-based classifier $U$, since $W^{con}$ is trained from scratch in a different space using ground-truth data and is hence not faithful to the original model, and 3) requires training a new $W^{con}$ classifier for every different set of concepts, hindering flexibility to the choice of concepts.
@@
Our method allows us to solve all the aforementioned issues and formulate zero-shot CBMs for any classifier. We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ represents the classification weights of the newly formulated classifier (the output of the text encoder $T$ for the class prompts). We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works (Rao et al., 2024), and without loss of generality, we use the $Z=20K$ most common words in English [^2] https://github.com/first20hours/google-10000-english as our concept set. These are general concepts that represent world knowledge and are not tailored towards any specific dataset. We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$. That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$. By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$. For an image $I$, we extract its visual features $f$ and use MLP to map them to $\tilde{f}$ which now lies in the text embedding space. That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$. Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to. That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and $C$. The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts. We provide an illustration in Figure 3.
Figure 3: Our proposed approach to concept discovery in CBMs.
@@
By observing Eq. 2, we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). Notably, if the gram matrix is identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, all we have to do is plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs are faithful to the original classifier. By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner, for any classifier, while also maintaining faithfulness to the original classifier and being flexible in the concept set.

### v2

# Unlocking Text Capabilities in Vision Models

@@ Abstract
Visual classifiers provide high-dimensional feature representations that are challenging to interpret and analyze. Text, in contrast, provides a more expressive and human-friendly interpretable medium for understanding and analyzing model behavior. We propose a simple, yet powerful method for reformulating any pretrained visual classifier so that it can be queried with free-form text without compromising its original performance. Our approach is label-free, data and compute-efficient, and is trained to preserve the underlying classifier’s distribution and decision-making processes. Our method unlocks several zero-shot text interpretability applications for any visual classifier. We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any visual classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language sentences. In both tasks we establish new state-of-the-art results, outperforming existing works and surpassing CLIP-based baselines with ImageNet-only trained classifiers, while using up to 400$\times$ fewer images and 400,000$\times$ less text during training.
@@ 1 Introduction
Visual classifiers provide dense, high-dimensional visual feature vectors that are difficult to interpret directly. Attribution maps—the standard way to interpret visual classifiers—only highlight the image region most influential to a prediction, offering a high-level spatial insight. They reveal where the key features lie, but not what those features actually mean or represent. By contrast, text is significantly more expressive, more precise and human-friendly, and it enables a new suite of text interpretability tasks  [1, 2, 3]. This raises a key question: How can we query a purely visual classifier with free-form text? Performing this allows us to have a text interface that can be used to query visual classifiers. Contrastive Language-Image Pre-Training (CLIP)-based vision–language models answer this by learning a shared embedding space [4], enabling any text to query the visual model directly. Yet in many real-world scenarios a high-performing, task-specific legacy model already exists and typically outperforms zero-shot vision-language models [5]. Retraining such a specialist on a large image–text corpus following the CLIP approach or the GPT-like approach is impractical, in terms of computational cost and need of a huge amount of image-text data. Moreover, this approach alters the decision-making process and distribution of the legacy model. As a result, most text-based interpretability techniques remain confined to CLIP-like models, leaving specialist vision models without a language-driven interface.
Figure 1: (a) Text interpretability is typically limited to models trained to share a joint embedding space such as CLIP-based ones. (b) Our approach removes this constraint, enabling any visual classifier to be queried through any arbitrary (unaligned) text encoder in a highly data- and compute-efficient, label-free manner, while preserving the classifier’s original distribution and reasoning process and thus offering text-based interpretability of the original classifier. The values here represent the cosine similarity between the text and the visual features $f$.
@@
In this work, we lift this limitation by proposing a method to reformulate any frozen visual classifier such that it can be accessed via free-form text (Figure 1). Our method is characterized by four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. The number of data points is also significantly reduced compared to CLIP-based approaches. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, it is explicitly trained to preserve the original distribution and reasoning process of the classifier. Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
@@
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. In ImageNet-1K [6], this set contains $1,000$ class labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are embedded into vector representations (e.g, using a text encoder or word embedding model), they provide semantic information. Specifically, these embeddings reside in a continuous space where nearby vectors capture related concepts. For the “goldfish" example, neighboring vectors might include “freshwater", “fins" and “orange". Our method learns to map images into this text embedding space using only class names, thus associating both the class name and its surrounding semantic space with the image. This is accomplished through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to output the same distribution scores across all classes, as the original classifier’s class distribution. This is done while while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with any text query beyond the class names.
@@
We demonstrate the effectiveness of our method with two primary applications. 1) We build label-free concept bottleneck models for any pretrained classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the newly built linear classifier shares the same reasoning process with the original classifier. 2) We show that our method can decode visual features of any classifier into natural language in a zero-shot manner, without requiring training on image-text data such as image-caption pairs, while also showing the generalization of our method to datasets beyond ImageNet.
In summary, our contributions are as follows: (i) We propose an efficient, label-free method to reformulate any visual classifier such that it can be queried with open-set text queries, maintaining its performance within a negligible margin (average of 0.2 points drop in accuracy) (ii) We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language. (iii) On both these applications, our method sets new state-of-the-art results outperforming existing works, including those using CLIP as the visual encoder, while being trained with 400$\times$ less images and 400,000$\times$ less text data.
@@ 2 Related Work
While CLIP is the predominant approach to interpret vision models through language [2], there exist works that try to decode visual features of models beyond CLIP. DeVIL [7] and LIMBER [8] trains an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data. Similarly, Natural Language Explanations (NLEs) [9, 10, 11] use annotated textual explanations in place of conventional captions. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process. ZS-A2T [12] converts attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective); therefore, it cannot be applied to standalone visual classifiers. In contrast, our method decodes visual features of any pretrained classifier in a zero-shot manner without requiring any annotated data, and is explicitly trained to keep the model’s predictive distribution intact.
Text-to-Concept (T2C) [13] and [14] are the closest related works to ours. These works train a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Since the linear layer maps features to the CLIP space, these works are strongly biased towards interpreting the CLIP model rather than the original classifier. As the ground-truth data for training the linear mapper are the features extracted from the CLIP vision encoder, this method is approximate to using the CLIP vision encoder directly to encode the image. Moreover, these works rely on CLIP supervision during training. In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution.
@@ 3 Method
Figure 2: An overview of our method for reformulating visual classifiers. (a) The process of training the MLP mapping between vision and text space. A PyTorch-like pseudocode for (a) in included in Section B of the Appendix. (b) The process of inference with the adapted visual classifier. The text encoder acts as a linear classifier weight generator in our re-formulated classification process. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
@@
We propose to learn a simple lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$. Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method efficient. We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}" where {class} represents the class name in text format. This results in $K$ textual prompts, each of which is encoded with $T$: $u_{i}=T(l^{p}_{i})$, $\forall i=1,\dots,K$. Stacking all the encoded prompts, we get a matrix $U\in\mathbb{R}^{K\times m}$. Here, $U$ acts as weights of the classification layer for our approach. We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$. Equivalently, this can be performed as a single vector-matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class. In other words, $S$ represents the classification logits of our approach.
The most straightforward way to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions) and sacrificing text interpretability of the original classifier. Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original distribution $o$ through the cross-entropy loss. For a single sample, the loss is given by:
@@
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but distill the distribution of the original model to a reformulated way of classification. Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters. Eq. 1 shows that our approach is label-free and steers the MLP to be faithful to the original classifier $F$, since it is explicitly trained for that purpose. Note also that this loss function naturally encodes the classifier’s relationship across all classes.
Our training approach is illustrated in Figure 2(a), and we provide a PyTorch-like pseudocode of it in Section B of the Appendix. It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section I in the Appendix for more information).
After training, the projected visual features and the text encoder features lie in the same space, and we can therefore query the visual features with any text by finding the alignment score between the encoded text and the visual features. In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the newly formulated classification task, defined as argmax($\tilde{f}.U^{T}$). This is shown in Figure 2(b).
@@ 4.1 Zero-Shot Concept Bottleneck Models
Concept Bottleneck Models (CBMs) [1] are a class of inherently interpretable models and have recently attracted significant attention. They consist of two steps: (1) concept discovery, followed by (2) concepts-to-predictions. In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image. In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of opaque features. Recently, Label-Free CBMs (LF-CBMs) use CLIP to perform the concept discovery step without annotated image-concept data. This is achieved either by using CLIP as a model to provide ground-truth image-concept similarities to train a layer that performs concept discovery [15], or by querying the image features from a set of predefined concepts within the CLIP space [16, 17] and using the cosine similarity scores between the concepts and the image as concept activations. However, there are major limitations in this existing line of CBMs. Firstly, LF-CBMs are restricted to CLIP-based models, as the concept discovery step would otherwise not be possible since it necessitates an image-text similarity model. Secondly, all CBM approaches to date require training a linear probe on the concept activations, which 1) hinders an on-the-fly ready-to-deploy CBM at test-time directly, and 2) requires training a new $W^{con}$ classifier from scratch for every different set of concepts, hindering flexibility to the choice of concepts.
Our method solves all the aforementioned issues, formulating zero-shot CBMs for any pretrained classifier. We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier. We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works [18], and without loss of generality, we use the $Z=20K$ most common words in English [19] as our concept set. These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset. To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set. Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”). In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name. This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes.
@@
We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$. That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$. By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$. For an image $I$, we extract its visual features $f$ and use the MLP to map them to $\tilde{f}$ which now lies in the text embedding space. That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$. Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to. That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and each row-vector in $C$. The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts. We provide an illustration in Figure 3(a).
Figure 3: (a) Our approach to concept discovery in CBMs. (b) Decoding visual features into natural language using any pretrained language decoder (e.g., GPT-2). We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features. Shown for iteration $j=1$
@@
From Eq. 2 we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). Notably, if the gram matrix is identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, all we have to do is plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs preserve the basis reasoning process of the original classifier. By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner and for any classifier. Note that our method is flexible in the concept set. The concept set is chosen at test time directly, and the CBM is built directly on-the-fly.